# Fundamentals and Advances in Transverse Thermoelectrics

### v1

@@ II.2 Transverse thermoelectric effects with structural symmetry breaking
In an artificial multilayer structure constructed by alternately stacking two conductors and cutting the stack at a certain angle, anisotropic transport properties occurs even if the constituent conductors exhibit isotropic electron/hole transport. When the off-diagonal terms of the thermoelectric transport tensor become finite due to such anisotropy, the artificially tilted multilayer (ATML) functions as a transverse thermoelectric conversion element [22, 23, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100] [Fig. 3(g)]. Since this transverse thermoelectric conversion originates from the Seebeck coefficient, it is called the off-diagonal Seebeck effect, which is clearly different from the Nernst effects that appear under time-reversal symmetry breaking. As described later, the phenomenological formulation differs in part between the transverse thermoelectric conversion principles in systems with time-reversal symmetry breaking and structural symmetry breaking.
@@
The transverse thermoelectric conversion based on the off-diagonal Seebeck effect has been studied for a long time [22, 23], and it is possible to design a transverse thermopower and figure of merit by selecting appropriate constituent materials and optimizing their tilt angle and thickness ratio. While the figure of merit for the off-diagonal Seebeck effect in ATML is improved by a high transverse thermopower, high effective electrical conductivity, and low effective thermal conductivity for the hybrid material, the characteristics required of each constituent material differ from those of conventional thermoelectric materials. In addition to the large difference in the Seebeck coefficient between the two materials, the contrast in the electrical and thermal conductivities between them is important, which induces nonuniform charge and heat current distributions [Fig. 3(g)]. In other words, even if the difference in the Seebeck coefficient is large, if the electrical and thermal conductivities of the two materials are the same, the transverse thermoelectric conversion does not occur. The transverse thermopower takes its maximum value when the tilt angle is 45, but the other parameters exhibit different behaviors. The smaller the tilt angle, the higher the effective electrical conductivity in the direction of the charge current and the lower the thermal conductivity in the direction of the heat current in ATML. Based on these situations, the maximum value of the figure of merit for the off-diagonal Seebeck effect in ATML is obtained at an angle smaller than 45 [99, 100]. However, if the tilt angle is too small, it is difficult to process and synthesize ATML in practice and its thermoelectric performance deteriorates due to the influence of boundary conditions [90, 91]. Thus, the contrast in the electrical and thermal conductivities should not be too large. Many studies on ATMLs have been conducted using sintered bulk materials and $z_{xy}T$ values exceeding 0.2 have been achieved around room temperature [100]. The transverse thermoelectric conversion performance of ATML can be calculated analytically, but in many systems, the experimentally obtained $z_{xy}T$ is much smaller than the ideal value due to the interfacial electrical and thermal resistances between the two materials. Despite this situation, in SmCo5/Bi0.2Sb1.8Te3 (BST) ATML developed by Ando et al., extremely small interfacial electrical and thermal resistances were achieved, and $z_{xy}T$ close to the ideal value was obtained experimentally, highlighting the importance of interface engineering in ATML [100].
@@
Even without constructing ATMLs, the transverse thermoelectric conversion can be achieved in single crystals showing anisotropic electronic conduction properties. In the single crystals of several metals such as Sb, Gd, Ho, Er, and Bi, the Seebeck coefficient is of the same sign but anisotropic [101, 102, 103]. Such materials having “unipolar” anisotropy make it possible to construct a transverse thermoelectric conversion element simply by cutting them so that the off-diagonal components of the thermoelectric transport tensor are finite. It is difficult to construct high-output power generation or cooling devices from such single-crystal metals because of their small transverse thermopower caused by the unipolar anisotropy in the Seebeck coefficient. In fact, due to their unipolar nature, the temperature gradient can never be purely orthogonal to the electric field [104]. Nevertheless, the off-diagonal Seebeck effect in single-crystalline metals is being considered for use in heat flux sensors because of their low thermal resistance [105].
It is also possible for single crystals to exhibit ADCP, though early experimental reports [106, 107, 108, 109] did not recognize their utility as transverse thermoelectrics until it was pointed out [27]. In recent years, even more single-crystalline materials with ADCP have been discovered [24, 25, 26, 110, 111, 112]. Unlike the aforementioned unipolar anisotropy, such materials exhibit “ambipolar anisotropy”, in which the sign of the Seebeck coefficient reverses depending on the crystal orientation. Therefore, by cutting the material along the direction between the axis showing the large positive Seebeck coefficient and the axis showing the large negative Seebeck coefficient and by applying a temperature gradient to allow the conduction electrons and holes to diffuse in the transverse direction, excellent transverse thermoelectric performance can be obtained through the off-diagonal Seebeck effect [see Fig. 3(h)]. For example, it has been reported that $z_{xy}T$ of single-crystalline Re4Si7 reaches 0.7 at 980 K [25]. This value is outstanding compared to other principles, but there are issues such as the limited choice of materials, the need for large single crystals, and the fact that the large figure of merit is obtained only at high temperatures.
@@ III Formulation of transverse thermoelectric conversion with time-reversal symmetry breaking
In this section we discuss transverse devices with time-reversal symmetry breaking, while in the next section we discuss transverse devices with structural symmetry breaking. An important distinction between time-reversal symmetry breaking and structural symmetry breaking is that the off-diagonal components of the Seebeck, electronic thermal conductivity, and electrical resistivity tensors [see Eqs. (25)–(27) below] are antisymmetric ($S_{xy}=-S_{yx}$, $\kappa_{xy}=-\kappa_{yx}$, $\rho_{xy}=-\rho_{yx}$) for the former and symmetric ($S_{xy}=S_{yx}$, $\kappa_{xy}=\kappa_{yx}$, $\rho_{xy}=\rho_{yx}$) for the latter [126]. It is a useful distinction for clarifying how fundamentally different these phenomena are from a tensor representational standpoint.
@@ III.1 Revising performance of longitudinal thermoelectric generator
The determinant of $\widehat{\cal A}$ is given by
$\det\widehat{\cal A}=\kappa\rho(1+zT),$
where $\kappa=\kappa_{p}+\kappa_{n}$, $\rho=\rho_{p}+\rho_{n}$, and the quantity
@@
where $V_{\rm emf}$ is the electromotive force and $R_{\rm tot}=R+R_{\rm L}$ is the total resistance of the circuit. Using $V_{\rm emf}=S\Delta T$ and ${\cal Q}_{\rm h}=Q_{x}$, the efficiency can be expressed as a function of $u=R_{\rm L}/R$ as
$\overline{\eta}=\frac{u}{\frac{(1+u)^{2}}{zT}+1+u},$
where $zT$ is defined in Eq. (9). The maximum reduced efficiency $\overline{\eta}$ is obtained for the stationary value $u^{*}=\sqrt{1+zT}$, which yields
@@
The determinant of $\widehat{\cal B}$ is calculated to be
$\det\widehat{\cal B}=\frac{R}{K^{\prime}}(1-z^{\prime}T),$
where
@@
is a new figure of merit defined by using $\kappa^{\prime}$ [120]. Note that, for Eq. (18) to be physically stable, $\det\widehat{\cal B}$ should be positive, from which we obtain a condition $01-z^{\prime}T>$ 0. This means that the figure of merit is bounded as
$10\leq z^{\prime}T<$ 1 . (20)
We next calculate the reduced efficiency for the present choice of independent variables. Using ${\cal Q}_{\rm h}=Q_{x}$, $V_{\rm emf}=S\Delta T$, and $J_{x}=V_{\rm emf}/R_{\rm tot}$ with $R_{\rm tot}=R_{\rm L}+R$, the reduced efficiency [Eq. (12)] is calculated to be
$\overline{\eta}=\frac{u}{\frac{(1+u)^{2}}{z^{\prime}T}-(1+u)},$
where $u=R_{\rm L}/R$. The maximum reduced efficiency $\overline{\eta}$ is obtained for the stationary value $u^{*}=\sqrt{1-z^{\prime}T}$, which yields
@@ III.2 Performance of transverse thermoelectric generator
$\widehat{\cal C}=\begin{pmatrix}\frac{-TS^{\rm(a)}_{xy}}{{K}^{\rm(a)}(l_{x}/l_{y})}&\frac{1}{{K}^{\rm(a)}}\\ -{R}^{\rm(a)}&\frac{-{S}^{\rm(a)}_{xy}}{{K}^{\rm(a)}(l_{x}/l_{y})}\end{pmatrix}.$ (40)
Note that in moving from Eqs. (37) and (38) to Eqs. (39) and (40), we used the symmetry $S_{yx}=-S_{xy}$ and $\kappa_{yx}=-\kappa_{xy}$. The determinant of $\widehat{\cal C}$ is calculated as
$\det\widehat{\cal C}=\frac{{R}^{\rm(a)}}{{K}^{\rm(a)}}(1+z_{xy}^{\rm(a)}T),$
where the quantity
@@
The remaining steps to calculate the reduced efficiency [Eq. (12)] of the transverse device in the adiabatic limit are basically the same as in the previous section. The thermal power input from the hot reservoir is given by ${\cal Q}_{\rm h}=Q_{x}$ and the current flowing through the circuit is given by $J_{y}=V_{\rm emf}/R_{\rm tot}$, where $R_{\rm tot}=R_{\rm L}+{R}^{\rm(a)}$ with $R_{\rm L}$ being the load resistance. Using $V_{\rm emf}=(-{S}^{\rm(a)}_{xy}(l_{y}/l_{x})/{K}^{\rm(a)}){\cal Q}_{\rm h}$, the reduced efficiency is calculated as
$\overline{\eta}=\frac{u}{\frac{(1+u)^{2}}{z_{xy}^{\rm(a)}T}+(1+u)},$
where $u=R_{\rm L}/R_{\rm tot}$. Then, the maximum reduced efficiency $\overline{\eta}$ is obtained for the stationary value $u^{*}=\sqrt{1+z_{xy}^{\rm(a)}T}$, which yields
@@
$\widehat{\cal D}=\begin{pmatrix}TS_{xy}(l_{y}/l_{x})&K\\ -R&-S_{xy}(l_{y}/l_{x})\end{pmatrix}.$ (49)
@@
$z_{xy}^{\rm(i)}T=\frac{TS_{xy}^{2}}{\rho_{yy}\kappa_{xx}}$ (50)
is the isothermal figure of merit for the transverse device. Note that, for Eq. (48) to be physically stable, $\det\widehat{\cal D}$ should be positive. Then, we have a condition $01-z_{xy}^{\rm(i)}T>$ 0, which means that $z_{xy}^{\rm(i)}T$ is bounded, namely,
$10\leq z_{xy}^{\rm(i)}T<$ 1 . (51)
Finally, we calculate the reduced efficiency $\overline{\eta}$ [Eq. (12)]. Using $V_{\rm emf}=S_{xy}(l_{y}/l_{x})\Delta T$ and $R_{\rm tot}=R+R_{\rm L}$ in Eq. (13), $\overline{\eta}$ can be calculated as
$\overline{\eta}=\frac{u}{\frac{(1+u)^{2}}{z_{xy}^{\rm(i)}T}-(1+u)},$
where $u=R_{\rm L}/R$. The maximum efficiency is obtained for the stationary value $u^{*}=\sqrt{1-z_{xy}^{\rm(i)}T}$, which yields
@@
$\kappa^{\prime\prime}_{xx}=\kappa_{xx}-\frac{TS_{xy}^{2}}{\rho_{yy}}$ (53)
@@
where $K^{\prime\prime}=\kappa^{\prime\prime}_{xx}l_{y}l_{z}/l_{x}$ and $R=\rho_{yy}l_{x}l_{z}/l_{y}$. Then, using $V_{\rm emf}=-S_{xy}(l_{y}/l_{x})\Delta T$ and $J_{y}=V_{\rm emf}/R_{\rm tot}$, where $R_{\rm tot}=R_{\rm L}+{R}^{\rm(a)}$, the reduced efficiency [Eq. (12)] is calculated to be
$\overline{\eta}=\frac{u}{\frac{(1+u)^{2}}{z^{\prime}_{xy}T}+1+u},$
where $u=R/R_{\rm L}$, and the corresponding figure of merit is given by
$z^{\prime\prime}_{xy}T=\frac{S^{2}_{xy}}{\rho_{yy}\kappa^{\prime\prime}_{xx}},$ (56)
@@
To summarize this subsection, we have formulated the maximum efficiency of a transverse thermoelectric generator for two thermal boundary conditions in the transverse direction: the adiabatic boundary condition [Eq. (43)], and isothermal boundary condition [Eq. (52)]. We note here that, for the Seebeck-driven anomalous/ordinary Hall effects and spin-Seebeck/spin-dependent-Seebeck-driven ISHEs, the efficiency expression is different from what we have discussed in the previous subsections, because these effects involve composite/hybrid materials [66, 83]. For example, the efficiency for the SSE device depends on the thickness ratio of the two layers of a magnetic insulator and attached metal.

### v2

@@ II.2 Transverse thermoelectric effects with structural symmetry breaking
In an artificial multilayer structure constructed by alternately stacking two conductors and cutting the stack at a certain angle, anisotropic transport properties occurs even if the constituent conductors exhibit isotropic electron/hole transport. When the off-diagonal components of the thermopower tensor become finite due to such anisotropy, the artificially tilted multilayer (ATML) functions as a transverse thermoelectric conversion element [22, 23, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109] [Fig. 3(g)]. Since this transverse thermoelectric conversion originates from the Seebeck coefficient, it is called the off-diagonal Seebeck effect, which is clearly different from the Nernst effects that appear under time-reversal symmetry breaking. As described later, the phenomenological formulation differs in part between the transverse thermoelectric conversion principles in systems with time-reversal symmetry breaking and structural symmetry breaking.
@@
The transverse thermoelectric conversion based on the off-diagonal Seebeck effect has been studied for a long time [22, 23], and it is possible to design a transverse thermopower and figure of merit by selecting appropriate constituent materials and optimizing their tilt angle and thickness ratio. While the figure of merit for the off-diagonal Seebeck effect in ATML is improved by a high transverse thermopower, high effective electrical conductivity, and low effective thermal conductivity for the hybrid material, the characteristics required of each constituent material differ from those of conventional thermoelectric materials. In addition to the large difference in the Seebeck coefficient between the two materials, the contrast in the electrical and thermal conductivities between them is important, which induces nonuniform charge and heat current distributions [Fig. 3(g)]. In other words, even if the difference in the Seebeck coefficient is large, if the electrical and thermal conductivities of the two materials are the same, the transverse thermoelectric conversion does not occur. The transverse thermopower takes its maximum value when the tilt angle is 45, but the other parameters exhibit different behaviors. The smaller the tilt angle, the higher the effective electrical conductivity in the direction of the charge current and the lower the thermal conductivity in the direction of the heat current in ATML. Based on these situations, the maximum value of the figure of merit for the off-diagonal Seebeck effect in ATML is obtained at an angle smaller than 45 [107, 108]. However, if the tilt angle is too small, it is difficult to process and synthesize ATML in practice and its thermoelectric performance deteriorates due to the influence of boundary conditions [98, 99]. Thus, the contrast in the electrical and thermal conductivities should not be too large. Many studies on ATMLs have been conducted using sintered bulk materials and $z_{xy}T$ values exceeding 0.2 have been achieved around room temperature [108]. The transverse thermoelectric conversion performance of ATML can be calculated analytically, but in many systems, the experimentally obtained $z_{xy}T$ is much smaller than the ideal value due to the interfacial electrical and thermal resistances between the two materials. Despite this situation, in SmCo5/Bi0.2Sb1.8Te3 (BST) ATML developed by Ando et al., extremely small interfacial electrical and thermal resistances were achieved, and $z_{xy}T$ close to the ideal value was obtained experimentally, highlighting the importance of interface engineering in ATML [108]. The transverse thermoelectric conversion performance of SmCo5/BST ATML has been further improved by utilizing the anisotropic electrical and thermal conductivities of the SmCo5 magnets and hybridizing their ANE; $z_{xy}T$ now reaches 0.3 around room temperature [109].
@@
Even without constructing ATMLs, the transverse thermoelectric conversion can be achieved in single crystals showing anisotropic electronic conduction properties. In the single crystals of several metals such as Sb, Gd, Ho, Er, and Bi, the Seebeck coefficient is of the same sign but anisotropic [110, 111, 112]. Such materials having “unipolar” anisotropy make it possible to construct a transverse thermoelectric conversion element simply by cutting them so that the off-diagonal components of the thermopower tensor are finite. It is difficult to construct high-output power generation or cooling/heating devices from such single-crystal metals because of their small transverse thermopower caused by the unipolar anisotropy in the Seebeck coefficient. In fact, due to their unipolar nature, the temperature gradient can never be purely orthogonal to the electric field [113]. Nevertheless, the off-diagonal Seebeck effect in single-crystalline metals is being considered for use in heat flux sensors because of their low thermal resistance [114].
It is also possible for single crystals to exhibit ADCP, though early experimental reports [115, 116, 117, 118] did not recognize their utility as transverse thermoelectrics until it was pointed out [27]. In recent years, even more single-crystalline materials with ADCP have been discovered [24, 25, 26, 119, 120, 121, 122]. Unlike the aforementioned unipolar anisotropy, such materials exhibit “ambipolar anisotropy”, in which the sign of the Seebeck coefficient reverses depending on the crystal orientation. Therefore, by cutting the material along the direction between the axis showing the large positive Seebeck coefficient and the axis showing the large negative Seebeck coefficient and by applying a temperature gradient to allow the conduction electrons and holes to diffuse in the transverse direction, excellent transverse thermoelectric performance can be obtained through the off-diagonal Seebeck effect [see Fig. 3(h)]. For example, it has been reported that $z_{xy}T$ of single-crystalline Re4Si7 reaches 0.7 at 980 K [25]. This value is outstanding compared to other principles, but there are issues such as the limited choice of materials, the need for large single crystals, and the fact that the large figure of merit is obtained only at high temperatures.
@@ III Formulation of transverse thermoelectric conversion with time-reversal symmetry breaking
In this section we discuss transverse devices with time-reversal symmetry breaking, while in the next section we discuss transverse devices with structural symmetry breaking. An important distinction between time-reversal symmetry breaking and structural symmetry breaking is that the off-diagonal components of the thermopower, electronic thermal conductivity, and electrical resistivity tensors [see Eqs. (25)–(27) below] are antisymmetric ($S_{xy}=-S_{yx}$, $\kappa_{xy}=-\kappa_{yx}$, $\rho_{xy}=-\rho_{yx}$) for the former and symmetric ($S_{xy}=S_{yx}$, $\kappa_{xy}=\kappa_{yx}$, $\rho_{xy}=\rho_{yx}$) for the latter [136]. It is a useful distinction for clarifying how fundamentally different these phenomena are from a tensor representational standpoint.
@@ III.1 Revising performance of longitudinal thermoelectric generator
The determinant of $\widehat{\cal A}$ is given by $\det\widehat{\cal A}=\kappa\rho(1+zT)$, where $\kappa=\kappa_{p}+\kappa_{n}$, $\rho=\rho_{p}+\rho_{n}$, and the quantity
@@
where $V_{\rm emf}$ is the electromotive force and $R_{\rm tot}=R+R_{\rm L}$ is the total resistance of the circuit. Using $V_{\rm emf}=S\Delta T$ and ${\cal Q}_{\rm h}=Q_{x}$, the efficiency can be expressed as a function of $u=R_{\rm L}/R$ as $\overline{\eta}={u}\{{{(1+u)^{2}}/{zT}+1+u}\}^{-1}$, where $zT$ is defined in Eq. (9). The maximum reduced efficiency $\overline{\eta}$ is obtained for the stationary value $u^{*}=\sqrt{1+zT}$, which yields
@@
The determinant of $\widehat{\cal B}$ is calculated to be $\det\widehat{\cal B}=({R}/{K^{\prime}})(1-z^{\prime}T)$, where
@@
is a new figure of merit defined by using $\kappa^{\prime}$ [130]. Note that, for Eq. (18) to be physically stable, $\det\widehat{\cal B}$ should be positive, from which we obtain a condition $1-z^{\prime}T>0$. This means that the figure of merit is bounded as
$0\leq z^{\prime}T<1$ (20)
We next calculate the reduced efficiency for the present choice of independent variables. Using ${\cal Q}_{\rm h}=Q_{x}$, $V_{\rm emf}=S\Delta T$, and $J_{x}=V_{\rm emf}/R_{\rm tot}$ with $R_{\rm tot}=R_{\rm L}+R$, the reduced efficiency [Eq. (12)] is calculated to be $\overline{\eta}={u}\{{(1+u)^{2}}/{z^{\prime}T}-(1+u)\}^{-1}$, where $u=R_{\rm L}/R$. The maximum reduced efficiency $\overline{\eta}$ is obtained for the stationary value $u^{*}=\sqrt{1-z^{\prime}T}$, which yields
@@ III.2 Performance of transverse thermoelectric generator
$\widehat{\cal C}=\begin{pmatrix}\frac{TS^{\rm(a)}_{yx}}{{K}^{\rm(a)}(l_{x}/l_{y})}&\frac{1}{{K}^{\rm(a)}}\\ -{R}^{\rm(a)}&\frac{{S}^{\rm(a)}_{yx}}{{K}^{\rm(a)}(l_{x}/l_{y})}\end{pmatrix}.$ (40)
Note that in moving from Eqs. (37) and (38) to Eqs. (39) and (40), we used the symmetry $S_{yx}=-S_{xy}$ and $\kappa_{yx}=-\kappa_{xy}$. The determinant of $\widehat{\cal C}$ is calculated as $\det\widehat{\cal C}=({{R}^{\rm(a)}}/{{K}^{\rm(a)}})(1+z_{xy}^{\rm(a)}T)$, where the quantity
@@
The remaining steps to calculate the reduced efficiency [Eq. (12)] of the transverse device in the adiabatic limit are basically the same as in the previous section. The thermal power input from the hot reservoir is given by ${\cal Q}_{\rm h}=Q_{x}$ and the current flowing through the circuit is given by $J_{y}=V_{\rm emf}/R_{\rm tot}$, where $R_{\rm tot}=R_{\rm L}+{R}^{\rm(a)}$ with $R_{\rm L}$ being the load resistance. Using $V_{\rm emf}=(-{S}^{\rm(a)}_{xy}(l_{y}/l_{x})/{K}^{\rm(a)}){\cal Q}_{\rm h}$, the reduced efficiency is calculated as $\overline{\eta}={u}\{{(1+u)^{2}}/{z_{xy}^{\rm(a)}T}+(1+u)\}^{-1}$, where $u=R_{\rm L}/R_{\rm tot}$. Then, the maximum reduced efficiency $\overline{\eta}$ is obtained for the stationary value $u^{*}=\sqrt{1+z_{xy}^{\rm(a)}T}$, which yields
@@
$\widehat{\cal D}=\begin{pmatrix}-TS_{yx}(l_{y}/l_{x})&K\\ -R&S_{yx}(l_{y}/l_{x})\end{pmatrix}.$ (49)
@@
$z_{xy}^{\rm(i)}T=\frac{TS_{yx}^{2}}{\rho_{yy}\kappa_{xx}}$ (50)
is the isothermal figure of merit for the transverse device. Note that, for Eq. (48) to be physically stable, $\det\widehat{\cal D}$ should be positive. Then, we have a condition $1-z_{xy}^{\rm(i)}T>0$, which means that $z_{xy}^{\rm(i)}T$ is bounded, namely,
$0\leq z_{xy}^{\rm(i)}T<1$ (51)
Finally, we calculate the reduced efficiency $\overline{\eta}$ [Eq. (12)]. Using $V_{\rm emf}=S_{xy}(l_{y}/l_{x})\Delta T$ and $R_{\rm tot}=R+R_{\rm L}$ in Eq. (13), $\overline{\eta}$ can be calculated as $\overline{\eta}={u}\{(1+u)^{2}/{z_{xy}^{\rm(i)}T}-(1+u)\}^{-1}$, where $u=R_{\rm L}/R$. The maximum efficiency is obtained for the stationary value $u^{*}=\sqrt{1-z_{xy}^{\rm(i)}T}$, which yields
@@
$\kappa^{\prime\prime}_{xx}=\kappa_{xx}-\frac{TS_{yx}^{2}}{\rho_{yy}}$ (53)
@@
where $K^{\prime\prime}=\kappa^{\prime\prime}_{xx}l_{y}l_{z}/l_{x}$ and $R=\rho_{yy}l_{x}l_{z}/l_{y}$. Then, using $V_{\rm emf}=-S_{xy}(l_{y}/l_{x})\Delta T$ and $J_{y}=V_{\rm emf}/R_{\rm tot}$, where $R_{\rm tot}=R_{\rm L}+{R}^{\rm(a)}$, the reduced efficiency [Eq. (12)] is calculated to be $\overline{\eta}={u}\{{(1+u)^{2}}/{z^{\prime}_{xy}T}+1+u\}^{-1}$, where $u=R/R_{\rm L}$, and the corresponding figure of merit is given by
$z^{\prime\prime}_{xy}T=\frac{S^{2}_{yx}}{\rho_{yy}\kappa^{\prime\prime}_{xx}},$ (56)
@@
To summarize this subsection, we have formulated the maximum efficiency of a transverse thermoelectric generator for two thermal boundary conditions in the transverse direction: the adiabatic boundary condition [Eq. (43)], and isothermal boundary condition [Eq. (52)]. We note here that, for the Seebeck-effect-driven anomalous/ordinary Hall effects and SSE/SdSE-driven ISHEs, the efficiency expression is different from what we have discussed in the previous subsections, because these effects involve composite/hybrid materials [74, 91]. For example, the efficiency for the SSE device depends on the thickness ratio of the two layers of a magnetic insulator and attached metal.